OpenStudy (aravindg):

Let f(X) be continuous funcion dfined for 1<=x<=3 .if f(x) takes rational values for all values of x and f(2)=10 . Determine value of f(1.5) @Physics

OpenStudy (amistre64):

hmm

OpenStudy (aravindg):

............

OpenStudy (amistre64):

is this simliar to finding a constant of variation?

OpenStudy (aravindg):

i dunno

OpenStudy (amistre64):

it feels to be missing parts to me, but thats just a guess

OpenStudy (aravindg):

help

OpenStudy (amistre64):

there are in infinite number of continuous functions from 1 to 3 that could have f(2) = 10; at least in my mind

OpenStudy (amistre64):

f(x) = 10 f(2) = 10 f(1.5) = 10

OpenStudy (aravindg):

???

OpenStudy (amistre64):

i know right, thats how i feel. it almost seems like an initial value question

OpenStudy (aravindg):

how u got hat?

OpenStudy (amistre64):

simple, i defined a function f(x) = 10 that is continuous from 1 to 3 and found values that for x=2 and x=1.5

OpenStudy (aravindg):

y it cant be any othr value????

OpenStudy (amistre64):

it can be anything it wants to be; it just so happens that this is how i defined the function.... given no other parameters it fits

OpenStudy (amistre64):

f(x) = 3x+4 is another one that will work fine f(2) = 10 f(1.5) = 8.5

OpenStudy (aravindg):

oh so there is no fixed answer?

OpenStudy (amistre64):

none that i can see with the given information.

OpenStudy (amistre64):

any none vertical line is continuous from 1 to 3; any parabola should work as long as you aint gonna tilt it on an odd axis

OpenStudy (amistre64):

cubics seem good

OpenStudy (aravindg):

i hav some more questions plz help

OpenStudy (aravindg):

k?

OpenStudy (amistre64):

i cant read your mind; you actually have to print something out :)

OpenStudy (aravindg):

k

OpenStudy (aravindg):

let f be a real fn satisfying f(x+y+z)=f(x)f(y)f(z) for all real x ,y ,zif f(2)=4 and f'(0)=3 find f(0) and f'(2)

OpenStudy (amistre64):

then f(x+y+z) - f(x)f(y)f(z) = 0 might be a good start

OpenStudy (aravindg):

so

OpenStudy (amistre64):

f(2) = 4 would suggest to me that x+y+z=4 since the function doesnt seem to be defined to well at the moment

OpenStudy (amistre64):

err; rather x+y+z = 2

OpenStudy (aravindg):

.....

OpenStudy (amistre64):

perhaps we might accomodate this: [f(x+y+z) - f(x)f(y)f(z)]' = f'(x+y+z) (x'+y'+z') - ( f'(x)x' f(y)f(z)+f(x)f'(y)y' f(z)+f(x)f(y)f'(z)z' ) but thats a bit confusing to me

OpenStudy (amistre64):

youd prolly need zarkon to sort thru it

OpenStudy (anonymous):

Just a note.... Your answer for the initial question is correct, but there aren't an infinite number of functions that work. The question specifies that f takes only rational values, and the real line is dense with irrationals, which means that between any two rational numbers there's an infinite number of irrationals. It follows that if the function is continuous, and it doesn't take on any irrational values, it must be constant.

OpenStudy (amistre64):

yay!! my guess was correct, somehow :)